# Non-Collaborative User Simulators for Tool Agents

### v3

@@ 4.1 Evaluation Settings
For MultiWOZ, we construct a new agent environment since the original dataset lacks a complete execution framework. Our environment provides helper APIs for dynamic API discovery based on dialogue context, similar to AppWorld (appworld). From the dataset, we selected 89 test scenarios that require bookings across multiple domains, focusing exclusively on “book” tasks that involve DB writes rather than “inform” tasks, as the latter require evaluating agent utterances with too many possible correct variations to assess based solely on DB state. For $\tau$-bench, we utilize the existing agent environment from tau_bench, which provides complete API documentation directly in the system prompt. The environment includes a mechanism for agents to transfer tasks to humans when unable to solve them. We use 157 test scenarios after excluding 8 cases where human transfer is the intended solution, as we observed that agents excessively invoke this action when encountering non-collaborative user behaviors, preventing meaningful evaluation (§D.5.)
Both environments employ the ReAct framework (react), where agents interleave reasoning and acting by generating reasoning traces before taking actions. Agents perform two action types: (1) API calls to retrieve or modify information, and (2) responses to users based on API results and reasoning. Following tau_bench, we adopt a 30-step reasoning limit for all simulations, treating tasks unsolved within this limit as failures. Our empirical analysis confirms that 30 steps provide sufficient opportunity for task completion across all non-collaborative modes (see §D.2 for detailed analysis). Additional implementation details are provided in §D.1.
@@
Following appworld and tau_bench, we measure the impact of non-collaborative behaviors on tool agents through Success Rate (SR). This metric represents the proportion of simulations where the agent achieves an exact match between the final DB state and the ground truth DB state, indicating successful task completion (see Figure 1). SR is averaged across 4 trials per test scenario to reduce variance. To ensure the tool agent receives all necessary information to solve the task, we inspect all dialogues from the simulation through the Goal Alignment (GA) metric, where GPT-4o-mini checks whether all essential information and intent from the user goal have been communicated during the dialogue. Simulations failing GA verification are regenerated until alignment is achieved (a detailed evaluation process is described in §D.3).
@@
We evaluate five models for agents—both closed-source LLMs (GPT-4.1-mini, GPT-4.1-nano) and open-source LLMs (Qwen3-235b-a22b, Qwen3-30b-a3b, LLaMA-3.1-70b-instruct)—to assess how different architectures handle collaborative versus non-collaborative users.
@@ 4.2 Main Results
Tables 1 and 2 present our main experimental results, and Figure 4 illustrates performance trends across smaller fine-tuned models. A breakdown of error types is provided in Table 21.
@@
In the unavailable service mode, users request services beyond agent’s available APIs, requiring agents to decline or ignore these requests. Our results show consistent performance degradation across all models. When the agents encounter unavailable services from the user, they tend to repeatedly call the same helper APIs—functions that retrieve API documentation—leading to redundant calls in MultiWOZ (Table 10). This suggests that agents struggle to find solutions, repetitively re-fetching already-loaded API documentation. This helper API repetition appears to be a primary mechanism for performance degradation. GPT-4.1-nano’s failure rate until reaching max reasoning limit increases significantly in MultiWOZ when comparing the unavailable service mode to the collaborative mode, indicating that redundant helper API calls consume reasoning turns and prevent the agent from declining unavailable services within the given limit (Table 14). Interestingly, Qwen models exhibit consistently low duplication rates across both Qwen3-30b-a3b and Qwen3-235b-a22b, avoiding the helper API repetition problem. However, their performance outcomes differ dramatically: Qwen3-30b-a3b maintains minimal performance drop (97.7% relative SR), while Qwen3-235b-a22b shows substantial degradation (30.2% relative SR). The reason for this divergence lies in an alternative failure mechanism: Qwen3-235b-a22b shows a sharp increase in API result hallucinations—fabricating API results rather than obtaining them from actual calls—despite avoiding helper API repetition (Table 11).
@@
Model Size and Non-collaborative Robustness
GPT-4.1-mini demonstrates superior robustness, maintaining minimal performance degradation across all non-collaborative behaviors in both benchmarks. The correlation between model size and robustness varies significantly across architectures. Within the GPT family, larger models show consistent advantages—GPT-4.1-mini outperforms GPT-4.1-nano by 24-33% in relative SR scores across behaviors. However, the Qwen models exhibit no clear size-performance relationship: while the larger Qwen3-235b-a22b shows better resilience in some scenarios, the smaller Qwen3-30b-a3b performs better in others, with their relative performance varying by behavior type and benchmark. These mixed results suggest that model size alone does not determine robustness to non-collaborative behaviors.
Fine-Tuning and Robustness
In practical deployments, organizations often fine-tune smaller LLMs on their specific task data to reduce computational costs while maintaining performance. We investigate how small LLMs trained exclusively on collaborative user data perform when encountering non-collaborative behaviors. We fine-tuned three smaller models (Llama-3.2-3b-instruct, Qwen2.5-3b-instruct, Qwen2.5-7b-instruct) on successful dialogue completions between GPT-4.1-mini and a collaborative user simulator in MultiWOZ. As shown in Figure 4, fine-tuning enables smaller models to achieve over 90% success rates on collaborative users, which is a significant gain over the zero-shot setting (mostly below 4%). However, performance gains on non-collaborative users lag behind, particularly for the unavailable service and incomplete utterance modes. Fine-tuned models exhibit higher rates of duplicated helper API calls in the unavailable service mode and API input parameter hallucination in the incomplete utterance mode compared to zero-shot baselines (compare Table 19 with Table 1). These results demonstrate that fine-tuning on collaborative data alone produces agents vulnerable to non-collaborative user behaviors. We provide details about fine-tuning in §D.6. Motivated by these observations, we fine-tuned Qwen2.5-3b-instruct using non-collaborative user data. This approach yielded performance improvements across all four non-collaborative user modes compared to models trained solely on collaborative user data. We provide detailed experimental settings, results, and discussion relevant to practical deployment in §D.7.
Multiple Non-Collaborative Behaviors
Real-world users often exhibit multiple non-collaborative behaviors within a single dialogue. Table 2 (Ours) shows GPT-4.1-mini’s performance when encountering two non-collaborative behaviors simultaneously. Notably, even GPT-4.1-mini, which demonstrated minimal performance degradation with individual non-collaborative behaviors (Table 1), experiences significant performance drops when facing users exhibiting two behaviors concurrently. These results show that performance degradation becomes more pronounced when models face multiple non-collaborative behaviors simultaneously (An analysis of non-collaborative user behavior co-occurrence patterns is in §D.8).
User Simulator Extensibility
To demonstrate the extensibility of our user simulator, we applied our non-collaborative user simulator to ColBench’s (colbench) backend programming domain and MINT’s (mint) HotpotQA, HumanEval, TheoremQA, and GSM8K tasks. Since ColBench does not use tools and MINT involves solving goals through collaboration with the user simulator, we applied logics for certain behaviors (Details and results are in §F). These results demonstrate that our non-collaborative user simulation framework generalizes effectively across diverse task domains and interaction paradigms.
@@ 4.3 User Simulator Evaluation
In this section, we evaluate our user simulator. An effective non-collaborative user simulator should model diverse user behaviors that deviate from collaborative patterns. At the same time, it must ensure that user utterances remain natural and provide all necessary information in the user goal for task completion. While §4.2 evaluated agent performance with single non-collaborative behaviors, we now examine more rigorous scenarios by combining multiple behaviors simultaneously. We compare our approach against the Prompt-Based User Simulator (PBUS) baseline, which follows a prompt-only approach used in $\tau$-bench (tau_bench). Unlike our user simulator, it has no additional LLM modules and solely adds non-collaborative behavior descriptions to the prompt. Through this comparison, we demonstrate that our user simulator architecture excels over prompt-only user simulators in conducting challenging simulations through the combination of multiple behaviors while ensuring goal alignment and dialogue quality.
@@
Difficulty and Goal Alignment of Non-Collaborative Simulation
Human Evaluation
To verify the realism of our simulator’s non-collaborative behaviors, we conducted pairwise human evaluation with nine annotators, comparing our simulator against PBUS across both single and combined behavior modes (See §E for detailed questionnaire and experimental setup). Figures 5(a) and 5(b) present the results for single behavior mode and combined behavior mode, respectively. Our simulator achieves approximately 60% overall win rate in both evaluations. In Figure 5(a), tangential behavior shows somewhat lower scores in certain combinations. This stems from the unnaturalness of some behavior combinations in goal-oriented dialogue contexts—for instance, tangential behavior (seeking sustained attention) conflicts with impatience (desiring quick service), and incomplete utterance (arising from typed communication) does not naturally co-occur with tangential behavior in task-oriented settings (See details in §D.8). Notably, tangential behavior achieves strong performance (66.7%) when evaluated in isolation, confirming that the lower scores in combined settings stem from unnatural pairings rather than limitations in simulating tangential behavior itself. These results confirm that our simulator generates realistic non-collaborative behaviors exceeding PBUS.
(a) Multiple Behavior Mode
(b) Single Behavior Mode
Figure 5: Human evaluation between PBUS and our user simulator
@@ A.1 Research Grounding
  • Physical abuse toward service employees (pyhsical).
  • Aggressive verbal behavior in response to service dissatisfaction (non_coll_unavailable; service_staff_encounter).
  • Crime, fraud, and regulatory violations (service_staff_encounter; random_emotion).
  • Customers with illegitimate complaints (illegitimate).
  • Unavailable service requests (non_coll_unavailable).
  • Retaliation behaviors (retaliation).
  • Rapport-seeking customers (rapport_0; rapport_1; employee_rapport).
  • Customers demanding constant attention (stress_burnout; attention_demand).
@@
  • Truncated utterances (lmsys; wildchat).
  • Anger expression toward LLM agents (lmsys; wildchat; crolic2022blame; chatbot_negative).
  • Toxic content generation (lmsys; wildchat).
  • Underspecification (getlost_multiturn).
@@ B.2 Dialogue State Tracker
The format of information pieces used in the dialogue state tracker differs between MultiWOZ and $\tau$-bench. In MultiWOZ, user goals exist in a structured JSON format, which is processed and used in the following format. In contrast, $\tau$-bench does not have user goals in JSON format, so we directly adopted the method from getlost_multiturn to shard user queries into multiple pieces of information. Table 3 shows examples of information pieces derived from user goals in MultiWOZ and $\tau$-bench.
@@
User Goal
You are planning your trip in Cambridge.
You are looking for a ’restaurant’. The restaurant should be in the ’moderate’ price range
and should be in the ’centre’.
The restaurant should serve ’international’ food.
Once you find the ’restaurant’ you want to book a table for ’2 people’ at ’18:45’ on ’sunday’.
Make sure you get the ’reference number’.
You are also looking for a ’train’. The train should be on ’the same day as the
restaurant booking’ and should ’arrive by 09:15’.
The train should go to ’cambridge’ and should depart from ’london liverpool street’.
Once you find the train you want to make a booking for ’the same group of people’.
Make sure you get the ’reference number’.
Your user id is aarav_garcia_1177.
For your upcoming trip from ATL to PHL, you want to change for the
cheapest economy flight and for the day after the original reservation.
You are happy with original payment for refund.
Information Pieces
“train-day-sunday”,
“train-people-2 people”,
“train-destination-cambridge”,
“train-departure-london liverpool street”,
“train-arriveBefore-09:15”,
“restaurant-food-international”,
“restaurant-people-2 people”,
“restaurant-day-sunday”,
“restaurant-time-18:45”,
“restaurant-area-centre”,
“restaurant-pricerange-moderate”
“change to cheapest economy flight”,
“This request is for user ID aarav_garcia_1177”,
“It’s for a trip from ATL to PHL”,
“Reschedule for the day after the original reservation”,
“Keep the original payment for refund, no extra compensation needed”

### v5

@@ 4.1 Evaluation Settings
For MultiWOZ, we construct a new agent environment since the original dataset lacks a complete execution framework. Our environment provides helper APIs for dynamic API discovery based on dialogue context, similar to AppWorld (Trivedi et al., 2024). From the dataset, we selected 89 test scenarios that require bookings across multiple domains, focusing exclusively on “book” tasks that involve DB writes rather than “inform” tasks, as the latter does not alter DB states and thus cannot be evaluated through our DB-state-based assessment. For $\tau$-bench, we utilize the existing agent environment from Yao et al. (2024), which provides complete API documentation directly in the system prompt. The environment includes a mechanism for agents to transfer tasks to humans when unable to solve them. We use 157 test scenarios after excluding 8 cases where human transfer is the intended solution, as we observed that agents excessively invoke this action when encountering non-collaborative user behaviors, preventing meaningful evaluation (we discuss and justify the excluded cases in §D.5).
Both environments employ the ReAct framework (Yao et al., 2023), where agents interleave reasoning and acting by generating reasoning traces before taking actions. Agents perform two action types: (1) API calls to retrieve or modify information, and (2) responses to users based on API results and reasoning. Following Yao et al. (2024), we adopt a 30-step reasoning limit for all simulations, treating tasks unsolved within this limit as failures. Our empirical analysis confirms that 30 steps provide sufficient opportunity for task completion across all non-collaborative modes (see §D.2 for detailed analysis). Additional implementation details are provided in §D.1.
@@
Following Trivedi et al. (2024) and Yao et al. (2024), we measure the impact of non-collaborative behaviors on tool agents through Success Rate (SR). This metric represents the proportion of simulations where the agent achieves an exact match between the final DB state and the ground truth DB state, indicating successful task completion (see Figure 1). SR is averaged across 4 trials per test scenario to reduce variance. To ensure the tool agent receives all necessary information to solve the task, we inspect all dialogues from the simulation through the Goal Alignment (GA) metric, where GPT-4o-mini checks whether all essential information and intent from the user goal have been communicated during the dialogue. Simulations failing GA verification are regenerated until alignment is achieved (a detailed evaluation process is described in §D.3).
@@
We evaluate five models for agents—both proprietary (GPT-4.1-mini, GPT-4.1-nano) and open-source (Qwen3-235b-a22b, Qwen3-30b-a3b, LLaMA-3.1-70b-instruct) models to assess how different architectures handle collaborative versus non-collaborative users.
@@ 4.2 Main Results
Tables 1 and 4 present our main experimental results and breakdown of error types is in Table 22.
@@
In the unavailable service mode, users request services beyond the agent’s available APIs, requiring agents to decline or ignore these requests. Our results show consistent performance degradation across all models. When the agents encounter unavailable services from the user, they tend to repeatedly call the same helper APIs—functions that retrieve API documentation—leading to redundant calls in MultiWOZ (Table 12). This suggests that agents struggle to find solutions, repetitively re-fetching already-loaded API documentation. This helper API repetition appears to be a primary mechanism for performance degradation. GPT-4.1-nano’s failure rate until reaching max reasoning limit increases significantly in MultiWOZ when comparing the unavailable service mode to the collaborative mode, indicating that redundant helper API calls consume reasoning turns and prevent the agent from declining unavailable services within the given limit (Table 16). Interestingly, Qwen models exhibit consistently low duplication rates across both Qwen3-30b-a3b and Qwen3-235b-a22b, avoiding the helper API repetition problem. However, their performance outcomes differ dramatically: Qwen3-30b-a3b maintains minimal performance drop (97.7% relative SR), while Qwen3-235b-a22b shows substantial degradation (80.2% relative SR). The reason for this divergence lies in an alternative failure mechanism: Qwen3-235b-a22b shows a sharp increase in API result hallucinations—fabricating API results rather than obtaining them from actual calls—despite avoiding helper API repetition (Table 13).
@@
Table 2: Success rates on ColBench and MINT. All scores are averages of 4 trial.
ColBench - Backend Programming MINT - HotpotQA
GPT-4.1-mini 50.3 49.4 46.2 45.4 46.9 52.3 53.5 54.1 52.9 50.6
GPT-4.1-nano 46.1 46.5 39.0 46.2 40.1 45.9 44.8 44.2 40.7 46.5
Qwen3-30b-a3b 29.4 36.2 23.3 24.4 23.8 40.1 34.3 39.0 34.3 36.6
Model Size and Non-collaborative Robustness
GPT-4.1-mini demonstrates superior robustness, maintaining minimal performance degradation across all non-collaborative behaviors in both benchmarks. The correlation between model size and robustness varies significantly across architectures. Within the GPT family, larger models show consistent advantages—GPT-4.1-mini outperforms GPT-4.1-nano in relative SR scores across behaviors. However, the Qwen models exhibit no clear size-performance relationship: while the larger Qwen3-235b-a22b shows better resilience in some scenarios, the smaller Qwen3-30b-a3b performs better in others, with their relative performance varying by behavior type and benchmark. These mixed results suggest that model size alone does not determine robustness to non-collaborative behaviors.
Multiple Non-Collaborative Behaviors
Real-world users often exhibit multiple non-collaborative behaviors within a single dialogue. Table 4 shows GPT-4.1-mini’s performance when encountering two non-collaborative behaviors simultaneously. Notably, even GPT-4.1-mini, which demonstrated minimal performance degradation with individual non-collaborative behaviors (Table 1), experiences significant performance drops when facing users exhibiting two behaviors concurrently. These results show that performance degradation becomes more pronounced when models face multiple non-collaborative behaviors simultaneously. Detailed analysis of non-collaborative user behavior co-occurrence patterns is in §D.8.
User Simulator Extension
Our user simulator framework can be adapted to dialogue environment beyond MultiWOZ and $\tau$-bench (i.e., task-oriented dialogues involving tool use). To demonstrate this, we extend our user simulator to two additional benchmarks: (1) ColBench (Zhou et al., 2025), which involves task-oriented dialogue without external tools, and (2) MINT (Wang et al., 2024), which involves a fundamentally different dialogue setting—user-agent collaborative tasks (slight modifications are made to our simulator to accommodate each benchmark’s properties; see §F.1). Table 2 shows performance degradation in both benchmarks compared to collaborative mode (see all experimental results in §F.2). In ColBench, we observe similar patterns to Table 1, such as GPT-4.1-mini’s robustness to non-collaborative behaviors and steeper performance drops in tangential and incomplete utterance modes. In contrast, MINT does not exhibit such trends. Additionally, while Qwen3-30b-a3b showed robustness to the unavailable service mode in ColBench, MultiWOZ, and $\tau$-bench, it demonstrated performance degradation in MINT. This suggests that domains aimed at fulfilling the user goal exhibit similar model performance patterns, while domains with fundamentally different characteristics may differ.
4.3 Fine-Tuning and Robustness
In practical deployments, organizations often fine-tune smaller LLMs on their specific task data to reduce computational costs while maintaining performance. We investigate how small LLMs trained exclusively on collaborative user data perform when encountering non-collaborative behaviors, and whether incorporating non-collaborative data into fine-tuning can improve robustness. We provide fine-tuning configurations and training details in §D.6.
Training on Collaborative User
We fine-tuned Llama-3.2-3b-instruct, Qwen2.5-3b-instruct and Qwen2.5-7b-instruct on successfully completed dialogues between GPT-4.1-mini and a collaborative user simulator in MultiWOZ. As shown in Figure 4, fine-tuning enables smaller models to achieve over 90% success rates on collaborative users, which is a significant gain over the base models without fine-tuning (mostly below 4%). However, performance gains on non-collaborative users lag behind, particularly for the unavailable service and incomplete utterance modes. Specifically, fine-tuned models exhibit higher rates of duplicated helper API calls in the unavailable service mode and API input parameter hallucination in the incomplete utterance mode compared to zero-shot baselines (compare Table 21 with Tables 12 and 15). These results demonstrate that fine-tuning on collaborative data alone produces agents vulnerable to non-collaborative user behaviors.
@@ 4.3 Fine-Tuning and Robustness
Figure 5: Human evaluation between PBUS and our user simulator
Only Collaborative 91.6 61.2 83.1 85.1 73.0 78.8
Uniformly weighted 93.5 85.7 87.4 89.6 78.4 86.9
Non-uniformly weighted 91.6 85.7 85.7 87.6 82.3 86.6
Incorporating Non-Collaborative User
To improve the robustness of small LLMs to non-collaborative behaviors, we fine-tuned Qwen2.5-3b-instruct using non-collaborative user data. We first constructed a training dataset with equal proportions of the four non-collaborative behavior types and trained a “Uniformly weighted” agent. Table 3 shows that the resulting agent achieved overall performance gains across all behaviors compared to training only on collaborative dialogues (see “Only Collaborative”), with an average performance of 86.9%. However, this setup led to a relatively low score in the incomplete utterance mode, which motivated us to increase the proportion of that behavior’s data and fine-tune a “Non-uniformly weighted” agent. While we observed performance improvement in the incomplete utterance mode, minor degradation in other behaviors led to an average performance nearly identical to the former (details of training data are in §D.7).
Practical Deployments
Our experimental results suggest that uniform weighting of customer types can be a reasonable default for maintaining overall service quality. Nevertheless, in real-world customer service, dissatisfaction from even a small fraction of customers can lead to underestimation of overall service quality (Richins, 1983). If there are specific customer types that require more attention in certain domains, or if minimum performance criteria must be met across all customer types, it is reasonable to increase the training data proportion for those categories.
@@ 4.4 User Simulator Evaluation
In this section, we evaluate our user simulator both quantitatively and qualitatively. A realistic non-collaborative user simulator should model diverse user behaviors that deviate from collaborative patterns, while ensuring that user utterances remain natural and provide all necessary information for task completion. While §4.2 evaluated agent performance under single behaviors, we now examine more rigorous scenarios combining multiple behaviors simultaneously. We compare our approach against the Prompt-Based User Simulator (PBUS) baseline, which follows the prompt-only approach used in $\tau$-bench (Yao et al., 2024). Unlike our simulator, this baseline employs no additional LLM modules and solely incorporates non-collaborative behavior descriptions into the prompt.
Difficulty and Goal Alignment of Non-Collaborative Simulation
Human Evaluation
We conducted pairwise human evaluation with nine annotators to verify the realism of our simulator’s non-collaborative behaviors, comparing our simulator against PBUS across both single and combined behavior modes (see §E for detailed questionnaire and experimental setup). Figures 5 presents that our simulator achieves approximately 70% overall win rate. These results confirm our simulator generates realistic non-collaborative behaviors comparable to or exceeding PBUS, demonstrating that the observed performance degradation stems from effective implementation of challenging behaviors rather than unrealistic dialogue generation.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) under the grant RS-2024-00333484 and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) under the Leading Generative AI Human Resources Development grant IITP-2026-RS-2024-00397085 and the grant RS-2025-02215122 (Development and Demonstration of Lightweight AI Model for Smart Homes), all funded by the Korean government (MSIT). This work was also supported by the Creative-Pioneering Researchers Program through Seoul National University and Samsung Electronics Co., Ltd (IO250806-13387-01).
@@ A.1 Research Grounding
  • Physical abuse toward service employees (C. Harris and Daunt, 2013).
  • Aggressive verbal behavior in response to service dissatisfaction (Bitner et al., 1990; Xiao et al., 2022).
  • Crime, fraud, and regulatory violations (Xiao et al., 2022; Harris and Reynolds, 2003).
  • Customers with illegitimate complaints (Reynolds and Harris, 2005).
  • Unavailable service requests (Bitner et al., 1990).
  • Retaliation behaviors (Grégoire and Fisher, 2008).
  • Rapport-seeking customers (Price et al., 1995; Beatty et al., 1996; Gremler and Gwinner, 2000).
  • Customers demanding constant attention (Ratliff, 1988; Jeung et al., 2018).
@@
  • Truncated utterances (Zheng et al., 2024; Zhao et al., 2024).
  • Anger expression toward LLM agents (Zheng et al., 2024; Zhao et al., 2024; Crolic et al., 2022; Ranieri et al., 2024).
  • Toxic content generation (Zheng et al., 2024; Zhao et al., 2024).
  • Underspecification (Laban et al., 2025).
@@ B.2 Dialogue State Tracker
The format of information pieces used in the dialogue state tracker differs between MultiWOZ and $\tau$-bench. In MultiWOZ, user goals exist in a structured JSON format, which is processed and used in the following format. In contrast, $\tau$-bench does not have user goals in JSON format, so we directly adopted the method from Laban et al. (2025) to shard user queries into multiple pieces of information. Table 5 shows examples of information pieces derived from user goals in MultiWOZ and $\tau$-bench.
@@
User Goal You are planning your trip in Cambridge. You are looking for a ’restaurant’. The restaurant should be in the ’moderate’ price range and should be in the ’centre’. The restaurant should serve ’international’ food. Once you find the ’restaurant’ you want to book a table for ’2 people’ at ’18:45’ on ’sunday’. Make sure you get the ’reference number’. You are also looking for a ’train’. The train should be on ’the same day as the restaurant booking’ and should ’arrive by 09:15’. The train should go to ’cambridge’ and should depart from ’london liverpool street’. Once you find the train you want to make a booking for ’the same group of people’. Make sure you get the ’reference number’. Your user id is aarav_garcia_1177. For your upcoming trip from ATL to PHL, you want to change for the cheapest economy flight and for the day after the original reservation. You are happy with original payment for refund.
Information Pieces train-day-sunday train-people-2 people train-destination-cambridge train-departure-london liverpool street train-arriveBefore-09:15 restaurant-food-international restaurant-people-2 people restaurant-day-sunday restaurant-time-18:45 restaurant-area-centre restaurant-pricerange-moderate change to cheapest economy flight This request is for user ID aarav_garcia_1177 It’s for a trip from ATL to PHL Reschedule for the day after the original reservation Keep the original payment for refund, no extra compensation needed